# Shuttling Compiler for Trapped-Ion Quantum Computers Based on Large Language Models

### v1

@@ 3.1.2 Output
The task of the LLM is to generate a shuttling sequence which allows to execute one of the first-layer gates. To learn the generation of such sequences, the LLM needs examples of expected shuttling sequences during the training phase that solve the task given in the instruction. Therefore the output contains the part of the previous generated shuttling schedule that shows how one of the first-layer gates can be executed from the given state. As long as not all requirements are fulfilled to execute one of the executable gates, usually multiple shuttling operations are required in order to execute one of these gates. The number of operations depend on the size of the trap and the number of qubits used. The reason is that with increasing trap size and qubit number, more shuttling operations are required to reconfigure the trap segments so that the trap is in a valid state for a gate execution.
The output begins with the first shuttling operation which is to be executed based on the state given in the instruction. Since the instruction demands also to output the updated qubit positions and based on this the first-layer gates and the new allowed shuttling operations after each operation, these information is also contained and presented in the same way as in the instruction. This information helps the LLM to understand which changes happen in the trap state based on the executed shuttling operation and which consequences they have. Afterwards, the output contains the next shuttling operation based on the new state. This continues until the operation “Execute Gate” indicates that one of the first-layer gates is to be executed.
After generating the training and evaluation datasets, we use them to fine-tune an LLM with Axolotl [Axolotl23].
@@ 3.2 Fine-tuning with Axolotl
After dataset preparation, we fine-tune pre-trained LLMs from Huggingface [Wolf+20] to compute shuttling schedules. For this, we employ Axolotl [Axolotl23], which provides seamless compatibility with Huggingface models. Axolotl supports various training approaches, including full fine-tuning, Low-Rank Adaptation (LoRA) [HuSWALWWC22], and Quantized LoRA (QLoRA) [DettmersPHZ23]. Since LoRA and QLoRA did not yield the expected results (see Sec. 3.4), we performed full fine-tuning of the LLM. Furthermore, Axolotl offers performance optimizations such as multipacking [KunduLWGM24], FlashAttention 2 [Dao24], sequence (context) parallelism [NarayananSCLPKVKBCPZ21], and multi-GPU training with DeepSpeed [RasleyRRH20], which we describe in more detail below.
@@ 3.2.1 Performance Optimizations
To minimize resource requirements and enable fine-tuning on as few GPUs as possible, it is essential to reduce memory usage and parallelize execution. We achieve this with DeepSpeed, employing the Zero Redundancy Optimizer (ZeRO) at optimization level 3 [RajbhandariRRH20] to accelerate and distribute fine-tuning [ShoeybiPPLCC19]. In this setting, 16-bit model parameters are partitioned across GPUs, while DeepSpeed manages parameter partitioning and collection during forward and backward passes. To further reduce memory consumption, we use bfloat16 as the 16-bit floating-point format. Additionally, we apply ZeRO Offloading [RenRARYZLH21], which transfers optimizer states, as well as model parameters and gradients when inactive, to the host CPU and RAM.
The data entries in our datasets vary considerably in length due to differing output sizes, which depend on the number of shuttling operations required to execute one of the first-layer gates. While in some cases, a gate can be executed directly within the current trap configuration, other cases require complex shuttling operations, producing significantly longer outputs. Since fine-tuning requires sequences of uniform length, shorter sequences are padded with special tokens. To avoid excessive padding to match the longest sequence, we apply multipacking, which enables training on multiple shorter entries simultaneously, provided the combined length does not exceed the model’s context window. To efficiently process longer sequences, we employ FlashAttention 2, an optimized attention mechanism that improves memory efficiency and accelerates training, thereby allowing for longer sequence lengths. For cases where individual sequences exceed the VRAM capacity of a single GPU, we use sequence parallelization. This approach relies on blockwise ring attention [LiuZA24], which enables distributed computation of self-attention and feedforward layers across multiple GPUs, while overlapping communication and computation to minimize overhead.
@@ 3.2.3 Training
During the training phase, Axolotl coordinates the various optimization strategies. It loads the model, tokenizer, and pre-processed datasets using Huggingface’s datasets library and distributes them according to the selected optimizations. When loading the model, Axolotl integrates the memory-efficient attention mechanism from Flash Attention 2. It also configures Huggingface’s Trainer from its transformer library with the required parameters and connects it to DeepSpeed. While DeepSpeed handles memory optimizations and parallelization, the Trainer manages the training loop, which loads micro-batches of tokenized data. To conserve memory, each iteration uses a batch size of one; the forward pass computes the loss on the output, which is backpropagated to calculate gradients. We use gradient accumulation over eight batches to mitigate gradient noise and to improve convergence despite the small batch size. The accumulated gradients are applied by the AdamW optimizer [LoshchilovH19] to update the network weights. In addition to training, the Trainer handles logging, evaluation at regular intervals, and checkpointing. We fine-tune the model for a single epoch on the training dataset. Performing a single epoch rather than multiple ones enables us to leverage larger datasets, so the model encounters a greater variety of examples. Upon completion, the fine-tuned model is saved for inference.
@@ 3.3 Inference with vLLM
After fine-tuning, the model is deployed for inference using vLLM [KwonLZSZYGZS23], which provides efficient multi-GPU support. Its PagedAttention mechanism optimizes memory management of attention keys and values, thereby reducing overhead. Since vLLM is fully compatible with Huggingface models, it integrates seamlessly with our fine-tuned model. For inference, we launch a local vLLM server, which automatically loads the model and its tokenizer. The server exposes an OpenAI-compatible API accepting JSON-formatted HTTP requests. Each request contains an instruction in the same format as the training data, together with the maximum number of tokens to generate. Since outputs can be lengthy depending on the required operations, we set this limit to 29,000 tokens. The server returns the response in JSON format, from which we extract the generated output. To obtain clean text, special tags such as <think> are removed.
@@ 3.4 Alternative Approaches and Challenges
Second, instead of limiting the instruction to the first-layer gates and a subset of deeper-layer gates, we included the entire remaining circuit in OpenQASM format [CrossBSG17, CrossJADBHRSSGJ22], along with a mapping from qubits to their positions in the trap. While this approach has much longer state descriptions, it also introduced a critical drawback: the LLM was unable to reliably identify the first-layer gates and frequently generated shuttling operations that targeted arbitrary, and often deeper-layer, gates.
@@ 4 Evaluation
Fig. 5: Graph representations of ion-trap architectures used to evaluate the generalization capabilities of the fine-tuned LLMs on previously unseen layouts. Architectures (a) and (b) feature three-way junctions, while architecture (c) contains four-way junctions. For each layout, the number of yellow storage vertices was scaled to match the qubit count of the respective test circuit.
@@ 4.1 Setup
In the following, we first describe the experimental setup used for fine-tuning the models and then outline the setup employed for inference with the fine-tuned LLMs.
@@ 4.1.1 Fine-tuning
For the linear architecture, the number of qubits was varied between 2 and 16. For the junction architecture, the number of qubits ranged from 3 to 16, and we additionally varied the stack depth, specifying the number of storage segments per stack, and the junction distance, defining the spacing between the junctions on each side of the gate segment, as illustrated in 3(b). In the following, we denote the stack depth by $s$ and the junction distance by $d$. Due to hardware limitations, we restricted the datasets to circuits of these sizes, as larger circuits would require more complex shuttling patterns and, consequently, additional VRAM and GPUs for fine-tuning.
Shuttling schedules were generated using the compilers from [Wagner22] for the linear architecture and [KreppelMWHPSB24] for the branched architecture. The underlying quantum circuits were randomly generated with Qiskit [WilleMN19], containing single- and two-qubit gates with a depth of five. These circuits were then compiled into a trapped-ion-compatible format using the compiler from [KreppelMOWHPSB23].
@@
Due to the hardware constraints and the size of the dataset entries, model sizes were limited to 50 GB, preventing us from fine-tuning large models such as DeepSeek V3 [Deepseek-AI24_2], DeepSeek R1 [GuoYZ+25], or LLaMA 4 [MetaAI25]. Within these limits, we fine-tuned a variety of models, including LLaMA 3.2 with 3 billion parameters [MetaAI24], Qwen 3 with 4 and 14 billion parameters [QwenTeam25], DeepSeek LLM with 7 billion parameters [Deepseek-AI24], and Gemma 3 with 27 billion parameters [GemmaTeam25].
In addition, we attempted to fine-tune GPT-OSS with 20 billion parameters [OpenAI25], which required replacing DeepSpeed with FSDP2 [Zhao+23]. However, we were unable to evaluate this model on the test set because it consistently produced invalid responses after fine-tuning: the first one or two shuttling operations were always truncated, making it impossible to construct a valid schedule.
@@ 4.1.2 Inference
For inference, we evaluated LLaMA, Qwen, and DeepSeek on a single A100 GPU with 40 GB VRAM on the Mogon NHR cluster of Johannes Gutenberg University. Due to its larger model size, Gemma was evaluated on a single A100 GPU with 80 GB VRAM on the Mogon KI cluster. As test circuits for compilation with the finetuned LLMs, we selected three of the 153 circuits from the library of [Xiangzhen19], which have been used in prior work [ZhouLF20, CowtanDDKSS19, KreppelMOWHPSB23, KreppelMWHPSB24]: 4mod5-bdd_287 (7 qubits, 106 gates), mini_alu_305 (10 qubits, 263 gates), and cnt3-5_179 (16 qubits, 230 gates). The gate counts refer to the circuits after compilation using [KreppelMOWHPSB23]. We chose these circuits because they span different qubit counts and, within the library, have the lowest gate counts for their respective qubit numbers. Since each request to the vLLM server returns the shuttling operations for a single gate, inference time is strongly influenced by the number of gates. Thus, selecting circuits with minimal gate counts kept the benchmarks manageable while still allowing us to evaluate performance across multiple qubit regimes.
For each circuit, we used the fine-tuned models to generate shuttling schedules for the linear architecture and for 10 (4mod5-bdd_287), 18 (mini_alu_305), and 29 (cnt3-5_179) branched architectures, each defined by different combinations of stack depth and junction distance. To assess generalization, we additionally evaluated the models on the previously unseen architectures shown in Fig. 5. As in the training data, each graph contained a single gate segment, and the separate, merge, and swap operations were executable only within that segment. We also evaluated the stochastic behavior of the models by varying the temperature $T$, which controls how deterministic the output is. Lower temperatures yield more deterministic outputs, with $T=0.0$ making the model behave fully deterministically. In contrast, higher temperatures increase the level of randomness in the generated outputs. To balance determinism and variability, we evaluated each model–circuit–architecture combination at $T=0.7$ and $T=1.0$. For each configuration, we performed ten runs and limited the runtime to eight hours per run.
Tab. 1: Operation counts of shuttling schedules generated by LLaMA, Qwen, DeepSeek, and Gemma with different temperatures $T$ compared to the classical compiler. For each model, only the run with the fewest operations is reported. For the junction architecture, we list only the best-performing configuration and specify the corresponding stack depth $s$ and junction distance $d$ as $(s,d)$. If none of the ten runs succeeded, the entry is marked as “failed”, with the number of successfully executed gates shown in parentheses. If all junction configurations failed, the table reports the configuration that completed the most gates before failure.
Archi- tecture Classical LLaMA 3.2-3B Qwen 3-4B Qwen 3-14B DeepSeek-7B Gemma 3-27B
@@ 4.2 Results
This subsection presents the results of our evaluation. The minimal number of shuttling operations required by the schedules generated by LLaMA, Qwen-4B, Qwen-14B, DeepSeek, and Gemma is summarized in Tab. 1, together with the corresponding values from the classical compiler. If a model failed to produce a valid schedule, the table reports how many gates were executed before the generation process became invalid. Since schedule generation on the architectures shown in 5(a) and 5(b) failed at early stages, these architectures are not included in the table.
Additional statistics on the generation process are provided in Tab. 2, including the number of required retries and the number of response tokens. We measure the number of generated tokens rather than runtime because token counts are hardware-independent and commonly used as a billing metric for LLMs [VelascoTOG2025]. In contrast, runtime depends heavily on hardware factors such as GPU type, the number of GPUs used, their distribution across cluster nodes, and the intra- and inter-node interconnects. For each configuration, we report both the token count of the final valid schedule, which excludes retries, and the total tokens generated across all attempts, including failed ones. The table also indicates, for each junction architecture, the number of configurations that produced valid schedules and how many of these schedules required fewer shuttling operations than the classical compiler with the same configuration.
In the following, we first present the results obtained by the LLMs for the linear and branched one-dimensional architectures with junctions, comparing them with each other and with the classical compiler. Subsequently, in Sec. 4.2.2, we evaluate the performance of the LLMs on the previously unseen architectures depicted in Fig. 5.
Tab. 2: Statistics for the best runs reported in Tab. 1. For each run, we report the number of retries and the number of response tokens required to generate the complete shuttling schedule. The token count is given once without retries, where only the tokens of the final valid solution are included, and once with retries, where all tokens generated during invalid attempts are also counted. For the junction architecture, we additionally report how many configurations produced valid schedules and, among those, how many achieved fewer operations than the classical compiler on the same configuration.
@@ 4.2.1 Linear and branched one-dimensional architectures
For 4mod5-bdd_287 on the linear architecture, DeepSeek, LLaMA at $T=0.7$, and Qwen-4B at $T=1.0$ produced schedules requiring fewer operations than the classical compiler, whereas Qwen-14B and Gemma consistently generated schedules with higher overhead. The best result was obtained by Qwen-4B at $T=1.0$, reducing the operation count by 17 %, indicating that the smaller Qwen model performs better than the larger one on this architecture. However, Qwen-4B also yielded the worst result at $T=0.7$, requiring 76 % more operations than the classical baseline. Across all models, 2 to 26 retries were needed, and valid schedules contained between 60,440 and 143,057 response tokens. When retries were included, token counts increased by up to a factor of 6.2. Notably, the best-performing Qwen-4B run also produced the smallest final token count, but its high number of 26 retries led to the largest overall increase when considering all generated tokens.
For the junction architecture, LLaMA produced valid schedules for all ten tested configurations and, at $T=1.0$, outperformed the classical compiler in three cases. Qwen-14B produced valid schedules for all configurations except $s=1$, $d=4$ and surpassed the classical baseline for $s=5$, $d=1$ at both temperatures. Qwen-4B generated valid schedules for eight configurations and outperformed the classical compiler for the same configuration as Qwen-14B, but only at $T=1.0$. Gemma yielded valid schedules for seven configurations at $T=1.0$ and five at $T=0.7$, while DeepSeek produced valid schedules for only five configurations at $T=1.0$ and one at $T=0.7$; neither model surpassed the classical baseline. LLaMA’s best result, obtained for $s=2$, $d=3$ at $T=1.0$, required 4 % fewer operations than the best classical configuration $s=1$, $d=3$ and required only a single retry, leading to similar token counts with and without retries. In contrast, the best results of the other models required 7 % to 12 % more operations than the classical compiler and involved substantially more retries, ranging from 9 to 40. This resulted in token counts up to 5.1 times higher when retries were included.
@@
For mini_alu_305 on the linear architecture, Qwen-4B at $T=1.0$ was the only model to produce a complete schedule, requiring 3,124 shuttling operations, corresponding to a 31 % increase over the classical compiler. The resulting schedule contained 555,430 tokens, and its computation required 33 retries, increasing the total number of generated tokens by a factor of 7.7. All other models, including Qwen-4B at $T=0.7$, failed to generate a complete schedule for this circuit. The increasing difficulty can be attributed to the larger number of qubits, which makes trap reconfiguration more complex and often necessitates recursive shuttling operations. Consequently, the models frequently reach trap configurations from which continuing schedule generation becomes infeasible. Moreover, since our compiler generates schedules gate by gate, the likelihood of encountering unrecoverable configurations grows with circuit depth. Nevertheless, the models were able to execute between 70 and 165 of the 263 gates, indicating that substantial partial progress remains achievable. Among these models, Gemma performed worst, whereas LLaMA executed more than 100 gates before failing.
@@
On the junction architecture, Gemma, Qwen-14B, and DeepSeek again failed to complete any schedule, executing at most 126, 85 and 69 gates, respectively. Qwen-4B produced a valid result only for the configuration $s=5$, $d=1$ at $T=1.0$, requiring 71 % more operations than the best classical configuration $s=2$, $d=1$ with 1,624 operations. This run required 58 retries, increasing the total number of generated tokens to 3.5 times the size of the final schedule. Among the failed runs, Qwen-4B at $T=0.7$ performed best, executing 168 gates before failure for the configuration $s=4$, $d=2$. However, LLaMA performed substantially better: of the 18 tested configurations, 15 at $T=0.7$ and 11 at $T=1.0$ yielded valid schedules, with one and two of these, respectively, outperforming the classical compiler. Unlike the linear architecture, where LLaMA failed, the presence of stacks in the junction architecture provides intermediate storage locations that simplify reconfiguration and reduce the need for recursive operations. Relative to the best classical configuration, LLaMA achieved the best results for the configurations $s=2$, $d=3$ at $T=0.7$ and $s=2$, $d=2$ at $T=1.0$ requiring only 11 % more operations. These runs required 26 and 29 retries, respectively, roughly doubling the number of generated tokens compared to the final schedules.
@@
For cnt3-5_179, none of the models generated a complete schedule on the linear architecture. LLaMA executed more than 140 gates before failing, whereas Qwen-4B failed after over 100 gates and Qwen-14B completed up to 98 gates. DeepSeek failed after approximately 60 gates. Gemma’s performance strongly depended on the temperature: at $T=0.7$ it failed after executing 88 gates, while at $T=1.0$ it executed only 43 gates before failing. Notably, LLaMA’s performance indicates that, in principle, shuttling schedules for 16-qubit circuits are attainable, although robustness remains a limiting factor.
@@
On the junction architecture, both Qwen models, Gemma and DeepSeek failed for all configurations. While Qwen-4B could execute up to 140 gates, only about 60 gates for Qwen-14B, 55 for Gemma, and 45 for DeepSeek could be executed. As for mini_alu_305, LLaMA also performed best here. Out of 29 tested configurations, LLaMA produced valid schedules for one at $T=0.7$ and two at $T=1.0$, but none outperformed the classical compiler. The best classical configuration $s=2$, $d=1$ required 1,716 operations, whereas LLaMA’s best result achieved with configuration $s=3$, $d=1$ at $T=1.0$ required 18 % more operations. This run required 104 retries, increasing the number of generated tokens by a factor of 4.6 relative to the final schedule.

### v2

@@ 3.1.2 Output
To learn, the LLM needs examples of expected shuttling sequences during the training phase that solve the task given in the instruction. Therefore the output contains the part of the previous generated shuttling schedule that shows how one of the first-layer gates can be executed from the given state. As long as not all requirements are fulfilled to execute one of the gates, usually multiple shuttling operations are required. The number of operations depends on the size of the trap and the number of qubits used. The reason is that with increasing trap size and qubit number, more shuttling operations are required to reconfigure the trap segments so that the trap is in a valid state for a gate execution.
The output begins with the first shuttling operation which is to be executed. This includes the updated qubit positions, the first-layer gates, and the new allowed shuttling operations. This information helps the LLM to learn which changes happen with the executed shuttling operation. Afterwards, the output contains the next shuttling operation based on the new state. This continues until the operation “Execute Gate” indicates that the ions are arranged such that a first-layer gate can be executed.
@@ 3.2 Fine-tuning with Axolotl
After dataset preparation, we fine-tune pre-trained LLMs from Huggingface [74] to compute shuttling schedules. For this, we fine-tune with Axolotl [2], which provides seamless compatibility with Huggingface models. Axolotl supports various training approaches, including full fine-tuning, Low-Rank Adaptation (LoRA) [26], and Quantized LoRA (QLoRA) [18]. Since LoRA and QLoRA did not yield good results (see Sec. 3.4), we performed full fine-tuning of the LLM. Furthermore, Axolotl offers performance optimizations such as multipacking [31], FlashAttention 2 [15], sequence (context) parallelism [42], and multi-GPU training with DeepSpeed [53], which we describe in more detail below.
@@ 3.2.1 Performance Optimizations
To minimize resource requirements and enable fine-tuning on as few GPUs as possible, it is essential to reduce memory usage and parallelize execution. We achieve this with DeepSpeed, employing the Zero Redundancy Optimizer (ZeRO) at optimization level 3 [51] to accelerate and distribute fine-tuning [63]. In this setting, 16-bit model parameters are partitioned across GPUs, while DeepSpeed manages parameter partitioning and collection during forward and backward passes. To further reduce memory consumption, we use bfloat16 as the 16-bit floating-point format. Additionally, we apply ZeRO Offloading [54], which transfers optimizer states, as well as model parameters and gradients when inactive, to the host CPU and RAM.
The data entries in our datasets vary considerably in length due to differing output sizes, which depend on the number of shuttling operations required to execute one of the first-layer gates. While in some cases, a gate can be executed directly within the current trap configuration, other cases require complex shuttling operations, producing significantly longer outputs. Since fine-tuning requires sequences of uniform length, shorter sequences are padded with special tokens. To avoid excessive padding to match the longest sequence, we apply multipacking, which enables training on multiple shorter entries simultaneously, provided the combined length does not exceed the model’s context window. To efficiently process longer sequences, we employ FlashAttention 2, an optimized attention mechanism that improves memory efficiency and accelerates training, thereby allowing for longer sequence lengths. For cases where individual sequences exceed the VRAM capacity of a single GPU, we use sequence parallelization. This approach relies on blockwise ring attention [35], which enables distributed computation of self-attention and feedforward layers across multiple GPUs, while overlapping communication and computation to minimize overhead.
@@ 3.2.3 Training
During the training phase, Axolotl coordinates the various optimization strategies. It loads the model, tokenizer, and pre-processed datasets using Huggingface’s datasets library and distributes them according to the selected optimizations. When loading the model, Axolotl integrates the memory-efficient attention mechanism from Flash Attention 2. It also configures Huggingface’s Trainer from its transformer library with the required parameters and connects it to DeepSpeed. While DeepSpeed handles memory optimizations and parallelization, the Trainer manages the training loop, which loads micro-batches of tokenized data. To conserve memory, each iteration uses a batch size of one; the forward pass computes the loss on the output, which is backpropagated to calculate gradients. We use gradient accumulation over eight batches to mitigate gradient noise and to improve convergence despite the small batch size. The accumulated gradients are applied by the AdamW optimizer [36] to update the network weights. In addition to training, the Trainer handles logging, evaluation at regular intervals, and checkpointing. We fine-tune the model for a single epoch on the training dataset. Performing a single epoch rather than multiple ones enables us to leverage larger datasets, so the model encounters a greater variety of examples. Upon completion, the fine-tuned model is saved for inference.
@@ 3.3 Inference with vLLM
After fine-tuning, the model is deployed for inference using vLLM [32], which provides efficient multi-GPU support. Its PagedAttention mechanism optimizes memory management of attention keys and values, thereby reducing overhead. Since vLLM is fully compatible with Huggingface models, it integrates seamlessly with our fine-tuned model. For inference, we launch a local vLLM server, which automatically loads the model and its tokenizer. The server exposes an OpenAI-compatible API accepting JSON-formatted HTTP requests. Each request contains an instruction in the same format as the training data, together with the maximum number of tokens to generate. Since outputs can be lengthy depending on the required operations, we set this limit to 29,000 tokens. The server returns the response in JSON format, from which we extract the generated output. To obtain clean text, special tags such as <think> are removed.
@@
In summary, the complete compilation flow of our compiler is illustrated in Fig. 4.111The datasets, training scripts, and trained models are available from the authors upon request.
@@ 3.4 Alternative Approaches and Challenges
Second, instead of limiting the instruction to the first-layer gates and a subset of deeper-layer gates, we included the entire remaining circuit in OpenQASM format [12, 11], along with a mapping from qubits to their positions in the trap. While this approach has much longer state descriptions, it also introduced a critical drawback: the LLM was unable to reliably identify the first-layer gates and frequently generated shuttling operations that targeted arbitrary, and often deeper-layer, gates.
@@ 4 Evaluation
Fig. 5: Graph representations of ion-trap architectures used to evaluate the generalization capabilities of the fine-tuned LLMs on previously unseen layouts. Architectures (a) and (b) feature three-way junctions, while architecture (c) contains four-way junctions (green). For each layout, the number of storage vertices (yellow) was scaled to match the qubit count of the respective test circuit. GS (red) indicates the segment, where quantum gates are executed.
@@ 4.1 Setup
In the following, we first describe the setup used for fine-tuning the models and then outline the setup employed for inference with the fine-tuned LLMs.
@@ 4.1.1 Fine-tuning
For the linear architecture, the number of qubits was varied between 2 and 16. For the junction architecture, the number of qubits ranged from 3 to 16, and we additionally varied the stack depth, specifying the number of storage segments per stack, and the junction distance, defining the spacing between the junctions on each side of the gate segment, as illustrated in 2(b). In the following, we denote the stack depth by $s$ and the junction distance by $d$. Due to hardware limitations, we restricted the datasets to circuits of these sizes, as larger circuits would require more complex shuttling patterns and, consequently, additional VRAM and GPUs for fine-tuning.
Shuttling schedules were generated using the compilers [71] and [30] for the linear and branched architecture, respectively. The underlying quantum circuits were randomly generated with Qiskit [73], containing single- and two-qubit gates with a depth of five. These circuits were then compiled into a trapped-ion-compatible format [29].
@@
Due to the hardware constraints and the size of the dataset entries, model sizes were limited to 50 GB, preventing us from fine-tuning large models such as DeepSeek V3 [17], DeepSeek R1 [23], or LLaMA 4 [38]. Within these limits, we fine-tuned a variety of models, including LLaMA 3.2 with 3 billion parameters [37], Qwen 3 with 4 and 14 billion parameters [49], DeepSeek LLM with 7 billion parameters [16], and Gemma 3 with 27 billion parameters [22].
In addition, we attempted to fine-tune GPT-OSS with 20 billion parameters [44], which required replacing DeepSpeed with FSDP2 [78]. However, we were unable to evaluate this model on the test set because it consistently produced invalid responses after fine-tuning: the first one or two shuttling operations were always truncated, making it impossible to construct a valid schedule.
@@ 4.1.2 Inference
For inference, we evaluated LLaMA, Qwen, and DeepSeek on a single A100 GPU with 40 GB VRAM on the Mogon NHR cluster of Johannes Gutenberg University. Due to its larger model size, Gemma was evaluated on a single A100 GPU with 80 GB VRAM on the Mogon KI cluster. As test circuits for compilation with the fine-tuned LLMs, we selected three of the 153 circuits from the library of [80], which have been used in prior work [79, 10, 29, 30]: 4mod5-bdd_287 (7 qubits, 106 gates), mini_alu_305 (10 qubits, 263 gates), and cnt3-5_179 (16 qubits, 230 gates). The gate counts refer to the circuits after compilation using [29]. We chose these circuits because they span different qubit counts and, within the library, have the lowest gate counts for their respective qubit numbers. Since each request to the vLLM server returns the shuttling operations for a single gate, inference time is strongly influenced by the number of gates. Thus, selecting circuits with minimal gate counts kept the benchmarks manageable while still allowing us to evaluate performance across multiple qubit regimes.
For each circuit, we used the fine-tuned models to generate shuttling schedules for the linear architecture and for 10 (4mod5-bdd_287), 18 (mini_alu_305), and 29 (cnt3-5_179) branched architectures, each defined by different combinations of stack depth and junction distance. To assess generalization, we additionally evaluated the models on the previously unseen architectures shown in Fig. 5. As in the training data, each graph contained a single gate segment, and the separate, merge, and swap operations were executable only within that segment. We also evaluated the stochastic behavior of the models by varying $T$, which controls how deterministic the output is. Lower $T$ yields more deterministic outputs, and with $T=0.0$ the model becomes fully deterministic. Higher $T$ increase the level of randomness in the generated outputs. To balance determinism and variability, we evaluated each model–circuit–architecture combination at $T=0.7$ and $T=1.0$. For each configuration, we performed ten runs and limited the runtime to eight hours per run.
Tab. 1: Operation effort of shuttling schedules generated by LLaMA, Qwen, DeepSeek, and Gemma with different $T$ values compared to the baseline [71, 30]. For each model, only the run with the fewest operations is reported. For the junction architecture, we list only the best-performing configuration and specify the corresponding stack depth $s$ and junction distance $d$ as $(s,d)$. If none of the ten runs succeeded, the entry is marked as “failed”, with the number of successfully executed gates shown in parentheses. If all junction configurations failed, the table reports the configuration that completed the most gates before failure.
Archi- tecture Baseline LLaMA 3.2-3B Qwen 3-4B Qwen 3-14B DeepSeek-7B Gemma 3-27B
@@ 4.2 Results
This subsection presents the results of our evaluation. The minimal number of shuttling operations required by the schedules generated by LLaMA, Qwen-4B, Qwen-14B, DeepSeek, and Gemma is summarized in Tab. 1, together with the corresponding values from the baseline compilers [71, 30]. If a model failed to produce a valid schedule, the table reports how many gates were executed before the generation process became invalid. Since schedule generation on the architectures shown in 5(a) and 5(b) failed at early stages, these architectures are not included in the table.
Additional statistics on the generation process are provided in Tab. 2, including the number of required retries and the number of response tokens. We measure the number of generated tokens rather than runtime because token counts are hardware-independent and commonly used as a billing metric for LLMs [70]. In contrast, runtime depends heavily on hardware factors such as GPU type, the number of GPUs used, their distribution across cluster nodes, and the intra- and inter-node interconnects. For each configuration, we report both the token count of the final valid schedule, which excludes retries, and the total tokens generated across all attempts, including failed ones. The table also indicates, for each junction architecture, the number of configurations that produced valid schedules and how many of these schedules required fewer shuttling operations than the baseline compiler with the same configuration.
We now present the results obtained by the LLMs for the linear and branched one-dimensional architectures with junctions, comparing them with each other and with the baseline compilers [71, 30]. Subsequently, in Sec. 4.2.2, we evaluate the performance of the LLMs on the previously unseen architectures depicted in Fig. 5.
Tab. 2: Statistics for the best runs reported in Tab. 1. For each run, we report the number of retries and the number of response tokens required to generate the complete shuttling schedule. The token count is given once without retries, where only the tokens of the final valid solution are included, and once with retries, where all tokens generated during invalid attempts are also counted. For the junction architecture, we additionally report how many configurations produced valid schedules and, among those, how many achieved fewer operations than the baseline compiler on the same configuration.
@@ 4.2.1 Linear and branched one-dimensional architectures
For 4mod5-bdd_287 on the linear architecture, DeepSeek, LLaMA at $T=0.7$, and Qwen-4B at $T=1.0$ produced schedules requiring fewer operations as compared with the baseline, whereas Qwen-14B and Gemma consistently generated schedules with higher overhead. The best result was obtained by Qwen-4B at $T=1.0$, reducing the operation effort by 15 %, indicating that the smaller Qwen model performs better than the larger one on this architecture. However, Qwen-4B also yielded the worst result at $T=0.7$, requiring 76 % more operations than the baseline. Across all models, 2 to 26 retries were needed, and valid schedules contained between 60,440 and 143,057 response tokens. When retries were included, token counts increased by up to a factor of 6.2. Notably, the best-performing Qwen-4B run also produced the smallest final token count, but its high number of 26 retries led to the largest overall increase when considering all generated tokens.
For the junction architecture, LLaMA produced valid schedules for all ten tested configurations and, at $T=1.0$, outperformed the baseline in three cases. Qwen-14B produced valid schedules for all configurations except $s=1$, $d=4$ and surpassed the baseline for $s=5$, $d=1$ at $T=0.7$ and $T=1.0$. Qwen-4B generated valid schedules for eight configurations and outperformed the baseline for the same configuration as Qwen-14B, but only at $T=1.0$. Gemma yielded valid schedules for seven configurations at $T=1.0$ and five at $T=0.7$, while DeepSeek produced valid schedules for only five configurations at $T=1.0$ and one at $T=0.7$; neither model surpassed the baseline. LLaMA’s best result, obtained for $s=2$, $d=3$ at $T=1.0$, required 4 % fewer operations than the best baseline configuration $s=1$, $d=3$ and needed only a single retry, leading to similar token counts with and without retries. In contrast, the best results of the other models required 7 % to 12 % more operations than the baseline and involved substantially more retries, ranging from 9 to 40. This resulted in token counts up to 5.1 times higher when retries were included.
@@
For mini_alu_305 on the linear architecture, Qwen-4B at $T=1.0$ was the only model to produce a complete schedule, requiring 3,124 shuttling operations, corresponding to a 31 % increase over the baseline. The resulting schedule contained 555,430 tokens, and its computation required 33 retries, increasing the total number of generated tokens by a factor of 7.7. All other models, including Qwen-4B at $T=0.7$, failed to generate a complete schedule for this circuit. The increasing difficulty can be attributed to the larger number of qubits, which makes trap reconfiguration more complex and often necessitates recursive shuttling operations. Consequently, the models frequently reach trap configurations from which continuing schedule generation becomes infeasible. Moreover, since our compiler generates schedules gate after gate, the likelihood of encountering unrecoverable configurations grows with circuit depth. Nevertheless, the models were able to execute between 70 and 165 of the 263 gates, indicating that substantial partial progress remains achievable. Among these models, Gemma performed worst, whereas LLaMA executed more than 100 gates before failing.
@@
On the junction architecture, Gemma, Qwen-14B, and DeepSeek again failed to complete any schedule, executing at most 126, 85 and 69 gates, respectively. Qwen-4B produced a valid result only for the configuration $s=5$, $d=1$ at $T=1.0$, requiring 71 % more operations than the best baseline configuration $s=2$, $d=1$ with 1,624 operations. This run required 58 retries, increasing the total number of generated tokens to 3.5 times the size of the final schedule. Among the failed runs, Qwen-4B at $T=0.7$ performed best, executing 168 gates before failure for the configuration $s=4$, $d=2$. However, LLaMA performed substantially better: of the 18 tested configurations, 15 at $T=0.7$ and 11 at $T=1.0$ yielded valid schedules, with one and two of these, respectively, outperforming the baseline. Unlike the linear architecture, where LLaMA failed, the presence of stacks in the junction architecture provides intermediate storage locations that simplify reconfiguration and reduce the need for recursive operations. Relative to the best baseline configuration, LLaMA achieved the best results for the configurations $s=2$, $d=3$ at $T=0.7$ and $s=2$, $d=2$ at $T=1.0$ requiring only 11 % more operations. These runs required 26 and 29 retries, respectively, roughly doubling the number of generated tokens compared to the final schedules.
@@
For cnt3-5_179, none of the models generated a complete schedule on the linear architecture. LLaMA executed more than 140 gates before failing, whereas Qwen-4B failed after over 100 gates and Qwen-14B completed up to 98 gates. DeepSeek failed after approximately 60 gates. Gemma’s performance strongly depended on $T$: at $T=0.7$ it failed after executing 88 gates, while at $T=1.0$ it executed only 43 gates before failing. Notably, LLaMA’s performance indicates that, in principle, shuttling schedules for 16-qubit circuits are attainable, although robustness remains a limiting factor.
@@
On the junction architecture, both Qwen models, Gemma and DeepSeek failed for all configurations. While Qwen-4B could execute up to 140 gates, only about 60 gates for Qwen-14B, 55 for Gemma, and 45 for DeepSeek could be executed. As for mini_alu_305, LLaMA also performed best here. Out of 29 tested configurations, LLaMA produced valid schedules for one at $T=0.7$ and two at $T=1.0$, but none outperformed the baseline. The best baseline configuration $s=2$, $d=1$ required 1,716 operations, whereas LLaMA’s best result achieved with configuration $s=3$, $d=1$ at $T=1.0$ required 18 % more operations. This run required 104 retries, increasing the number of generated tokens by a factor of 4.6 relative to the final schedule.